# Spectral Moment Formulae for $\hbox{GL}(3)\times\hbox{GL}(2)$ L-functions II: The Eisenstein Case

### v2

@@ 6.1. Spectral expansion and Eisenstein contribution
Let $\mathfrak{M}_{\alpha}^{(3)}(s;H)$ be defined in (1.6). For $\sigma>3/2$ and $H\in\mathcal{C}_{\eta}$, we have
$2\cdot\mathcal{I}\big(s;\,P(\,*\,;h),\,\widetilde{\Phi}^{\mathrm{ND}}\big)\ =\ \mathfrak{M}_{\alpha}^{(3)}(s;H).$ (6.2)
@@
This follows from a modification of [Kw23, Proposition 5.25] and Corollary 3.22. ∎
The cuspidal part of $\mathfrak{M}_{\alpha}^{(3)}(s;H)$ evidently admits an entire continuation. The Eisenstein part, initially defined on $\sigma>1+\epsilon_{0}/2$, also admits a continuation, though this is less immediate. Define
@@
The function $s\mapsto\mathcal{C}_{\alpha}(s;0;H)$ admits a meromorphic continuation to $\sigma>1-\eta+\epsilon_{0}$. On the vertical strip $\epsilon_{0}<\sigma<1-\epsilon_{0}$, the continuation is given by $\mathcal{R}_{\alpha}(s;H)+\mathcal{C}_{\alpha}(s;0;H)$.
@@
This is an adaptation of [Mo97, p. 118] and we sketch the argument for our setting. Firstly, suppose $1+\epsilon_{0}/2<\sigma<1+\epsilon_{0}$. Since $H\in\mathcal{C}_{\eta}$, we may shift the line of integration to $\operatorname{Re}\mu=\eta>40$, picking up the second set of poles in (6.4) together with the zeros of $\Lambda(1-2\mu)$. We have
@@
as the meromorphic continuation of $\mathcal{C}_{\alpha}(s;0;H)$ to the domain $\epsilon_{0}<\sigma<1-\epsilon_{0}$. A little calculation shows that $\mathcal{R}_{\alpha}(s;H)=-\sum_{i=1}^{3}\,\operatorname{Res}_{\mu=s-1-\alpha_{i}}\,+\,\sum_{i=1}^{3}\,\operatorname{Res}_{\mu=1-s+\alpha_{i}}$, and this completes the sketch. ∎
@@
The expression $\mathfrak{M}_{\alpha}^{(3)}(s;H)$ admits a meromorphic continuation to $\sigma>1-\eta+\epsilon_{0}$. On the strip $\epsilon_{0}<\sigma<1-\epsilon_{0}$, the continuation is given by $\mathfrak{M}_{\alpha}^{(3)}(s;H)+\mathcal{R}_{\alpha}(s;H)$.
@@ 6.2. Unfolding and rearrangement
where the right-hand side converges absolutely on $1+\epsilon_{0}<\sigma<4$; see [Kw23, Section 6B]. In what follows, we perform our analysis on suitable vertical strips within the half-plane $\sigma<4$. To avoid confusion with computations in the literature (e.g., regarding sign convention of parameters and normalization of Fourier coefficients), it is better to consider instead
$\mathcal{J}(s;\,h,\,\Phi)\ :=\ \int_{0}^{\infty}\int_{0}^{\infty}\ h(y_{1})(y_{0}^{2}y_{1})^{s-\frac{1}{2}}\mathcal{P}_{\psi}(\mathbf{y};\,\Phi^{\mathrm{ND}})\ d^{*}\mathbf{y}.$ (6.8)
As is apparent from [Kw23, eqs. (6.5), (6.7), (6.8)], we have
@@ 6.3. Degenerate terms—I
Suppose $1/2+\epsilon_{0}<\sigma<4$. Then the function $\mathcal{J}_{(0,1)}^{\min}(s;\,h,\,\Phi)$ is equal to
@@
The function $\mathcal{J}_{(0,1)}^{\min}(s;\,h,\,\Phi)$ admits a holomorphic continuation to the strip $\epsilon_{0}<\sigma<4$, except for a simple pole at $s=(1+\alpha_{3}^{w})/2$.
@@
Recall that $H:=h^{\#}\in\mathcal{C}_{\eta}$ and $\eta>40$. Using the bound $W_{(\alpha_{1}^{w}-\alpha_{2}^{w})/2}(y)\,\ll_{A}\,y^{-A}$ with $A>2\sigma-3/2+\epsilon_{0}$ and $\eta-1-\epsilon_{0}>|A-\sigma|$, one observes that $J_{w,\alpha}(s;h)$ converges absolutely whenever $\epsilon_{0}<\sigma<4$. With the analytic continuation of the $\zeta$-function, the right-hand side of (6.17) serves as a continuation for $\mathcal{J}_{(0,1)}^{\min}(s;\,h,\,\Phi)$ to $\epsilon_{0}<\sigma<4$. It is clearly holomorphic on $\epsilon_{0}<\sigma<4$, except for the simple pole coming from the factor $\zeta(2s+\alpha_{1}^{w}+\alpha_{2}^{w})$. This completes the proof. ∎
@@
$\ \cdot\ \int_{(0)}\,H(\mu)\,\prod_{\pm}\ \Gamma_{\mathbb{R}}(s\pm\mu+\alpha_{1}^{w})\Gamma_{\mathbb{R}}(s\pm\mu+\alpha_{2}^{w})\Gamma_{\mathbb{R}}(1-s\pm\mu-\alpha_{3}^{w})\ d^{W}\mu.$ (6.21)
@@ 6.4. Degenerate terms—II
The main task of this section is to evaluate $\mathcal{J}_{(*,0)}^{\min}(s;\,h,\,\Phi)$ defined in Proposition 6.6. Interestingly, the $\gamma$-sum in (6.15), which arises from the $\Gamma_{3}$ Fourier expansion of $\Phi$, will transform into the $\Gamma_{2}$ Eisenstein series, and produce three of the main terms for the fourth moment of the $\zeta$-function. This is also essential for the analytic continuation of (6.7). Such phenomena do not appear in the first two moments of $\mathrm{GL}(2)$ $L$-functions ([Mo92]).
@@
$\mathcal{J}_{(*,0)}^{\min}(s;\,h,\,\Phi)\ =\ \sum_{w=w_{3},w_{5},w_{\ell}}\ H(s+\alpha_{1}^{w})\,\frac{\Lambda(1+\alpha_{1}^{w}-\alpha_{2}^{w})\Lambda(1+\alpha_{1}^{w}-\alpha_{3}^{w})\Lambda(2s+\alpha_{1}^{w}+\alpha_{2}^{w})\Lambda(2s+\alpha_{1}^{w}+\alpha_{3}^{w})}{\Lambda(1+2(s+\alpha_{1}^{w}))}.$ (6.24)
Remark 6.11.
[…]
@@
$\mathcal{J}_{(*,0)}^{\min}(s;\,h,\,\Phi)\ =\ \sum_{w=w_{3},w_{5},w_{\ell}}\$$\ \left(\sum_{n_{1}=1}^{\infty}\ \frac{\zeta(w,\alpha,\psi_{(n_{1},0)})}{|n_{1}|^{1-\alpha_{3}^{w}+(2s-1)}}\right)$
@@
$2\,\int_{0}^{\infty}\,h(y_{1})\,\int_{0}^{1}\,E\left(z;s+\alpha_{1}^{w}\right)e(-u)\ du\ \frac{dy_{1}}{y_{1}^{2}}\ =\ 2\,\frac{H(s+\alpha_{1}^{w})}{\Lambda(1+2(s+\alpha_{1}^{w}))}\,.$ (6.30)
Now, the desired result follows from (6.4), (6.4), (6.29) and (6.30), and this completes the proof. ∎
@@
The function $\mathcal{J}_{(*,0)}^{\min}(s;\,h,\,\Phi)$ admits a holomorphic continuation to $\epsilon_{0}<\sigma<4$, except for the six simple poles at $s=(1+\alpha_{3}^{w})/2,\ (1+\alpha_{2}^{w})/2$ ($w=w_{3},w_{5},w_{\ell}$).
6.5. Diagonal and preparation for off-diagonal: $\mathcal{J}^{\mathrm{reg}}(s;\,h,\,\Phi)$
@@
Recall Proposition 6.6 on the decomposition of periods. From [Kw23, eq. (6.7)], we have
@@
where the term $\mathrm{OD}_{\alpha}(s)$ denotes the “off-diagonal”, i.e., terms with $a_{0}a_{1}\neq 0$ in (6.13). The first term of (6.33) can be evaluated by (3.40), (3.13) and (3.10). By [Kw23, eq. (6.8)], we have
@@
where the expression (6.34) converges absolutely whenever $1+\epsilon_{0}<\sigma<4$ and for $H\in\mathcal{C}_{\eta}$ ($\eta>40$). By [Kw23, Proposition 7.2] and (3.40), this can be written as (6.32). This completes the proof. ∎
@@ 6.6. Symmetries of the integral transform
The function $s\mapsto\left(\mathcal{F}_{\alpha}H\right)\left(2s-1,s\right)$ admits a holomorphic continuation to $\epsilon_{0}<\sigma<4$ except for three simple poles $s=(1-\alpha_{i})/2$, $i=1,2,3$. On $\epsilon_{0}<\sigma<1-\epsilon_{0}$, we have
@@
The argument also appears in [Kw23]; for completeness, we include a proof here. Suppose $1/2+\epsilon_{0}<\sigma<4$ and $s_{0}=2s-1$. In (4), the factor $\Gamma\left(\frac{1-u}{2}\right)$ in the denominator cancels with $\Gamma\left(s-\frac{s_{0}+u}{2}\right)$ in the numerator. This leads to
@@ 6.7. Analytic continuation and polar terms
The analytic continuation argument performed in [Kw23] is robust and carries over to the present case with minor modifications.
@@
Readers are invited to consult [Kw23, Section 9] for fuller details. Parallel to [Kw23, Section 9A] (‘Step 11), we shift the line of integration in (6.32) to $\operatorname{Re}s_{0}=2\epsilon_{0}$. This time, however, we pick up the residues of three extra simple poles at $s_{0}=1-\alpha_{i}$ ($i=1,2,3$), which results in
@@
This serves as a holomorphic continuation to $1/2+\epsilon_{0}<\sigma<4$, except for the three simple poles at $s=1-\alpha_{i}/2$ ($i=1,2,3$). Parallel to [Kw23, Section 9B] (‘Step $2$’), we restrict to the strip $1/2+\epsilon_{0}<\sigma<3/4$, and shift the line of integration from $\operatorname{Re}s_{0}=2\epsilon_{0}$ to $\operatorname{Re}s_{0}=1/2$, crossing the simple pole of $\zeta(2s-s_{0})$ with residue equal to 1-1. The function on the second line of (6.7) is holomorphic on the strip $1/4+\epsilon_{0}<\sigma<3/4$.
Section 9C of [Kw23] (‘Step 3’) carries over to the present context without change as it merely takes care of the necessary regularity (using Proposition 4.1 and the imposed assumptions on our class of test functions). Section 9D of [Kw23] (‘Step 4’) concerns the continuation of (6.44) and (6.45). Here, we instead adopt an explicit approach and the desired conclusion follows directly from Proposition 6.146.15. Section 9E of [Kw23] (‘Step 5’) is postponed to Corollary 6.17 and Section 6.8. This completes the proof of Proposition 6.16. ∎
@@
$\mathfrak{M}_{\alpha}^{(3)}(s;H)\ +\ \mathcal{D}_{\alpha}(s;H)\ +\ \mathcal{R}_{\alpha}(s;H)$ (6.49)
$\hskip 30.0pt\ =\ \prod_{i=1}^{3}\,\zeta(2s+\alpha_{i})\int_{(0)}\,H(\mu)\,\prod_{i=1}^{3}\,\prod_{\pm}\,\Gamma_{\mathbb{R}}(s\pm\mu-\alpha_{i})\ d^{W}\mu$ (6.50)
+ $\hskip 80.0pt\ +\ 4\sum_{w=w_{2},w_{4},w_{\ell}}\,\ \Lambda(1+\alpha_{2}^{w}-\alpha_{3}^{w})\Lambda(1+\alpha_{1}^{w}-\alpha_{3}^{w})\zeta(2s+\alpha_{1}^{w}+\alpha_{2}^{w})J_{w,\alpha}(s;h)$ (6.51)
@@
We complete the proof of Theorem 1.3 by demonstrating that the terms (6.50)–(6.53) agree with the CFKRS predictions for the $\mathrm{SO}(\text{even})$ symmetry described in Section 5, i.e., with $\mathcal{M}_{\alpha}^{i}(s;H)$ $(i=0,1,2,3)$ defined in Theorem 1.3. Observe that:
@@
    (0-swap). By $\alpha_{1}+\alpha_{2}+\alpha_{3}=0$, eq. (6.50) can clearly be written as $\mathcal{M}_{\alpha}^{0}(s;H)$ as defined by (1.11).
  2. (2)
    (11-swap). Putting Propositions 6.7 and 6.9 together, a pair of factors $\Gamma_{\mathbb{R}}(1+\alpha_{2}^{w}-\alpha_{3}^{w})$ (resp. $\Gamma_{\mathbb{R}}(1+\alpha_{1}^{w}-\alpha_{3}^{w})$) cancels, and it follows that eq. (6.51) is equal to $\mathcal{M}_{\alpha}^{1}(s;H)$ as defined by (1.12).
@@
    it follows that eq. (6.52) coincides with $\mathcal{M}_{\alpha}^{2}(s;H)$ as defined by (1.13).
@@
    (33-swap). Using Proposition 6.15 and (6.55) thrice, Eq. (6.53) coincides with $\mathcal{M}_{\alpha}^{3}(s;H)$ as defined by (1.14).
We set $s=1/2$. The matching with the predictions in Section 5.1 follows from the Weyl actions (1.4). The restriction $\alpha_{1}+\alpha_{2}+\alpha_{3}=0$ can be removed by analytic continuation in $\alpha_{i}$’s.

### v3

@@ 6.1. Spectral expansion and Eisenstein contribution
Let $\mathfrak{M}_{-\alpha}^{(3)}(s;H)$ be defined in (1.6). For $\sigma>3/2$ and $H\in\mathcal{C}_{\eta}$, we have
$2\cdot\mathcal{I}\big(s;\,P(\,*\,;h),\,\widetilde{\Phi}^{\mathrm{ND}}\big)\ =\ \mathfrak{M}_{-\alpha}^{(3)}(s;H).$ (6.2)
@@
This follows from a modification of [49, Proposition 5.25] and Corollary 3.22. ∎
The cuspidal part of $\mathfrak{M}_{-\alpha}^{(3)}(s;H)$ evidently admits an entire continuation. The Eisenstein part, initially defined on $\sigma>1+\epsilon_{0}/2$, also admits a continuation, though this is less immediate. Define
@@
The function $s\mapsto\mathcal{C}_{\alpha}(s;0;H)$ admits a meromorphic continuation to $\sigma>1-\eta+\epsilon_{0}$. On the vertical strip $\epsilon_{0}<\sigma<1-\epsilon_{0}$, the continuation is given by $\mathcal{R}_{\alpha}(1-s;H)+\mathcal{C}_{\alpha}(s;0;H)$.
@@
This is an adaptation of [60, p. 118] and we sketch the argument for our setting. Firstly, suppose $1+\epsilon_{0}/2<\sigma<1+\epsilon_{0}$. Since $H\in\mathcal{C}_{\eta}$, we may shift the line of integration to $\operatorname{Re}\mu=\eta>40$, picking up the second set of poles in (6.4) together with the zeros of $\Lambda(1-2\mu)$. We have
@@
as the meromorphic continuation of $\mathcal{C}_{\alpha}(s;0;H)$ to the domain $\epsilon_{0}<\sigma<1-\epsilon_{0}$. A little calculation shows that $\mathcal{R}_{\alpha}(1-s;H)=-\sum_{i=1}^{3}\,\operatorname{Res}_{\mu=s-1-\alpha_{i}}\,+\,\sum_{i=1}^{3}\,\operatorname{Res}_{\mu=1-s+\alpha_{i}}$, and this completes the sketch. ∎
@@
The expression $\mathfrak{M}_{-\alpha}^{(3)}(s;H)$ admits a meromorphic continuation to $\sigma>1-\eta+\epsilon_{0}$. On the strip $\epsilon_{0}<\sigma<1-\epsilon_{0}$, the continuation is given by $\mathfrak{M}_{-\alpha}^{(3)}(s;H)+\mathcal{R}_{\alpha}(1-s;H)$.
@@ 6.2. Unfolding and rearrangement
where the right-hand side converges absolutely on $1+\epsilon_{0}<\sigma<4$; see [49, Section 6B]. In what follows, we perform our analysis on suitable vertical strips within the half-plane $\sigma<4$.
To avoid confusion with computations in the literature (e.g., regarding the sign convention of parameters and normalization of Fourier coefficients), it is better to consider instead
$\mathcal{J}(s;\,h,\,\Phi)\ :=\ \int_{0}^{\infty}\int_{0}^{\infty}\ h(y_{1})(y_{0}^{2}y_{1})^{s-\frac{1}{2}}\mathcal{P}_{\psi}(\mathbf{y};\,\Phi^{\mathrm{ND}})\ d^{*}\mathbf{y}$ (6.8)
in the main calculations of Sections 6.36.4. As is apparent from [49, (6.5), (6.7), (6.8)], we have
@@ 6.3. Degenerate terms—I
Suppose $1/2+\epsilon_{0}<\sigma<4$. Then $\mathcal{J}_{(0,1)}^{\min}(s;h,\Phi)$ is equal to
@@
The function $\mathcal{J}_{(0,1)}^{\min}(s;\,h,\,\Phi)$ admits a holomorphic continuation to the strip $\epsilon_{0}<\sigma<4$, except for three simple poles at $s=(1+\alpha_{i})/2$, $i=1,2,3$.
@@
Recall that $H:=h^{\#}\in\mathcal{C}_{\eta}$ and $\eta>40$. Using the bound
$W_{(\alpha_{1}^{w}-\alpha_{2}^{w})/2}(y)\,\ll_{A}\,y^{-A}$
with $A>2\sigma-3/2+\epsilon_{0}$ and $\eta-1-\epsilon_{0}>|A-\sigma|$, observe that $J_{w,\alpha}(s;h)$ converges absolutely whenever $\epsilon_{0}<\sigma<4$. With the analytic continuation of the $\zeta$-function, the right-hand side of (6.17) serves as a continuation for $\mathcal{J}_{(0,1)}^{\min}(s;\,h,\,\Phi)$ to $\epsilon_{0}<\sigma<4$. It is clearly holomorphic on $\epsilon_{0}<\sigma<4$, except for the simple poles coming from the factor $\zeta(2s+\alpha_{1}^{w}+\alpha_{2}^{w})$ ($w\in\{w_{2},w_{4},w_{\ell}\}$). This completes the proof. ∎
@@
$\hskip 30.0pt\ \cdot\ \int_{(0)}\,H(\mu)\prod_{\pm}\,\Gamma_{\mathbb{R}}(1-s\pm\mu-\alpha_{3}^{w})\prod_{i=1}^{2}\,\Gamma_{\mathbb{R}}(s\pm\mu+\alpha_{i}^{w})\,d^{W}\mu.$ (6.21)
@@
Corollary 6.10.
Let $\mathcal{M}_{-\alpha}^{1}(s;H)$ be defined in (1.10). We have
$\mathcal{J}_{(0,1)}^{\min}(s;h,\widetilde{\Phi})\ =\ \frac{1}{2}\,\mathcal{M}_{-\alpha}^{1}(s;H).$ (6.24)
Proof.
Putting Propositions 6.7 and 6.9 together, we notice that a pair of factors $\Gamma_{\mathbb{R}}(1-\alpha_{2}^{w}+\alpha_{3}^{w})$ (resp. $\Gamma_{\mathbb{R}}(1-\alpha_{1}^{w}+\alpha_{3}^{w})$) cancels. Explicating the Weyl actions of $w_{2},w_{4},w_{\ell}$ on the parameters $(\alpha_{1},\alpha_{2},\alpha_{3})$ described in (3.3.3) and replacing $\Phi\to\widetilde{\Phi}$, the desired result follows. ∎
@@ 6.4. Degenerate terms—II
The main task of this section is to evaluate $\mathcal{J}_{(*,0)}^{\min}(s;\,h,\,\Phi)$ defined in Proposition 6.6. Interestingly, the $\gamma$-sum in (6.15), which arises from the $\Gamma_{3}$ Fourier expansion of $\Phi$, will transform into the $\Gamma_{2}$ Eisenstein series, and produce three of the main terms for the fourth moment of the $\zeta$-function. This is also essential for the analytic continuation of (6.7). Such phenomena do not appear in the first two moments of $\mathrm{GL}(2)$ $L$-functions ([58]).
@@
$\mathcal{J}_{(*,0)}^{\min}(s;\,h,\,\widetilde{\Phi})\ =\ \mathcal{R}_{-\alpha}(s;H).$ (6.25)
@@
$\mathcal{J}_{(*,0)}^{\min}(s;\,h,\,\Phi)\ =\ \sum_{w=w_{3},w_{5},w_{\ell}}\$$\ \Big(\sum_{n_{1}=1}^{\infty}\ \frac{\zeta(w,\alpha,\psi_{(n_{1},0)})}{|n_{1}|^{1-\alpha_{3}^{w}+(2s-1)}}\Big)$
@@
$2\,\int_{0}^{\infty}\,h(y_{1})\,\int_{0}^{1}\,E(z;s+\alpha_{1}^{w})e(-u)\ du\ \frac{dy_{1}}{y_{1}^{2}}\ =\ 2\,\frac{H(s+\alpha_{1}^{w})}{\Lambda(1+2(s+\alpha_{1}^{w}))}\,.$ (6.31)
Now, from (6.4), (6.4), (6.30) and (6.31), it follows that $\mathcal{J}_{(*,0)}^{\min}(s;\,h,\,\Phi)$ is given by
$\sum_{w=w_{3},w_{5},w_{\ell}}\ H(s+\alpha_{1}^{w})\,\frac{\Lambda(1+\alpha_{1}^{w}-\alpha_{2}^{w})\Lambda(1+\alpha_{1}^{w}-\alpha_{3}^{w})\Lambda(2s+\alpha_{1}^{w}+\alpha_{2}^{w})\Lambda(2s+\alpha_{1}^{w}+\alpha_{3}^{w})}{\Lambda(1+2(s+\alpha_{1}^{w}))}.$ (6.32)
Recall the Weyl actions of $w_{3},w_{5},w_{\ell}$ on the parameters $(\alpha_{1},\alpha_{2},\alpha_{3})$ in (3.3.3) and replace $\Phi\to\widetilde{\Phi}$, the desired result follows. ∎
@@
[…]
@@
Recall Proposition 6.6 on the decomposition of periods. From [49, eq. (6.7)], we have
@@
where the term $\mathrm{OD}_{\alpha}(s)$ denotes the “off-diagonal”, i.e., terms with $a_{0}a_{1}\neq 0$ in (6.13). The first term of (6.35) can be evaluated by (3.39), (3.13) and (3.10). By [49, eq. (6.8)], we have
@@
where the expression (6.36) converges absolutely whenever $1+\epsilon_{0}<\sigma<4$ and for $H\in\mathcal{C}_{\eta}$ ($\eta>40$). By [49, Proposition 7.2] and (3.39), this can be written as (6.34). This completes the proof. ∎
@@ 6.6. Symmetries of the integral transform
The function $s\mapsto\left(\mathcal{F}_{\alpha}H\right)\left(2s-1,s\right)$ admits a holomorphic continuation to $\epsilon_{0}<\sigma<4$ except for three simple poles $s=(1-\alpha_{i})/2$$(i=1,2,3)$. On $\epsilon_{0}<\sigma<1-\epsilon_{0}$, we have
@@
The argument also appears in [49]; for completeness, we include a proof here. Suppose $1/2+\epsilon_{0}<\sigma<4$ and $s_{0}=2s-1$. In (4), the factor $\Gamma\left(\frac{1-u}{2}\right)$ in the denominator cancels with $\Gamma\left(s-\frac{s_{0}+u}{2}\right)$ in the numerator. This leads to
@@ 6.7. Analytic continuation and polar terms
The analytic continuation argument performed in [49] is robust and carries over to the present case with minor modifications.
@@
Readers are invited to consult [49, Section 9] for fuller details. Parallel to [49, Section 9A] (‘Step 11), we shift the line of integration in (6.34) to $\operatorname{Re}s_{0}=2\epsilon_{0}$. This time, however, we pick up the residues of three extra simple poles at $s_{0}=1-\alpha_{i}$ ($i=1,2,3$), which results in
@@
This serves as a holomorphic continuation to $1/2+\epsilon_{0}<\sigma<4$, except for the three simple poles at $s=1-\alpha_{i}/2$ ($i=1,2,3$). Parallel to [49, Section 9B] (‘Step $2$’), we restrict to the strip $1/2+\epsilon_{0}<\sigma<3/4$, and shift the line of integration from $\operatorname{Re}s_{0}=2\epsilon_{0}$ to $\operatorname{Re}s_{0}=1/2$, crossing the simple pole of $\zeta(2s-s_{0})$ with residue equal to 1-1. The function on the second line of (6.7) is holomorphic on the strip $1/4+\epsilon_{0}<\sigma<3/4$.
Section 9C of [49] (‘Step 3’) carries over to the present context without change as it merely takes care of the necessary regularity (using Proposition 4.1 and the imposed assumptions on our class of test functions). Section 9D of [49] (‘Step 4’) concerns the continuation of (6.46) and (6.47). Here, we instead adopt an explicit approach and the desired conclusion follows from Proposition 6.146.15. This completes the proof.  [^12] The content of [49, Section 9E] (‘Step 5’) is postponed to Corollary 6.17 and Section 6.8.
@@
$\mathfrak{M}_{-\alpha}^{(3)}(s;H)\ +\ \mathcal{R}_{-\alpha}(s;H)\ +\ \mathcal{R}_{\alpha}(1-s;H)$ (6.51)
$\hskip 30.0pt\ =\ \prod_{i=1}^{3}\,\zeta(2s+\alpha_{i})\int_{(0)}\,H(\mu)\,\prod_{i=1}^{3}\,\prod_{\pm}\,\Gamma_{\mathbb{R}}(s\pm\mu-\alpha_{i})\ d^{W}\mu\ \ \ +\ \mathcal{M}_{-\alpha}^{1}(s;H)$ (6.52)
@@
We complete the proof of Theorem 1.3 by demonstrating that the totality of the terms (6.52), (6.53) and (6.54) agree with the CFKRS predictions for the $\mathrm{SO}(\text{even})$ symmetry described in Section 5, i.e., with $\sum_{0\leq i\leq 3}\,\mathcal{M}_{-\alpha}^{i}(s;H)$ defined in Theorem 1.3. Indeed, observe that:
@@
    (0 & 11-swap). By $\alpha_{1}+\alpha_{2}+\alpha_{3}=0$, the first term of (6.52) can be written as $\mathcal{M}_{-\alpha}^{0}(s;H)$ as defined by (1.9). The second term of (6.52) is already in the form of the $1$-swap prediction.
@@
    it follows that eq. (6.53) coincides with $\mathcal{M}_{-\alpha}^{2}(s;H)$ as defined by (1.11).
@@
    (33-swap). Using Proposition 6.15 and (6.56) thrice, observe that (6.54) coincides with $\mathcal{M}_{-\alpha}^{3}(s;H)$ as defined by (1.12).
We set $s=1/2$. The matching with the predictions in Section 5.1 follows from the Weyl actions (3.3.3). The restriction $\alpha_{1}+\alpha_{2}+\alpha_{3}=0$ can be removed by analytic continuation in $\alpha_{i}$’s.